1. Compute the current flowing through a copper wire of cross-sectional area 1.25Γ—10βˆ’6 π‘š2 assuming that the drift velocity is 2Γ—10βˆ’4 π‘š/𝑠 and the charge carrier density is 1.0Γ—1029/π‘š3.

2. (a) Calculate the resistance per unit length of a 22-gauge Nichrome wire of radius 0.321 mm.

(b) If a potential difference of 10.0 V is maintained across a 1.00 m length of the Nichrome wire, what is the current in the wire?

3. A metal wire of diameter 2 π‘šπ‘š and length 300 π‘š has a resistance of 1.642 Ξ© at 20 ℃ and 2.415 Ξ© at 150 ℃. Find the values of 𝛼0, 𝑅0 and 𝜌20.


1
Expert's answer
2021-05-03T10:31:38-0400

(1.) Here,

Cross- sectional area "A=1.25\times 10^{-6} m^2"

Drift velocity "v_d=2\times 10^{-4}\ m/s"

Carrier Charge Density "n=1.0\times 10^{29} \ \" \"m^3"

Using the formula,

Current flowing through a copper wire

"I=neAv_d"

"= (1.0\times 10^{29})\times(1.6\times10^{-19})\times(1.25\times10^{-6})\times(2\times10^{-4})"

"= 4\ Amperes"

Hence, Current flowing through a copper wire is 4 A



(2.) (a) The cross-sectional area of the wire is

"A=\pi r^2=\pi(0.321\times 10^{-3}m)^2=3.24\times 10^{-7} \ m^2"

and the resistivity of Nichrome is "1.5\times 10^{-6}\ \ \Omega\cdot m"

So, the resistivity per unit length

"\dfrac{R}{l}=\dfrac{\rho}{A}=\dfrac{1.5\times 10^{-6}\ \ \Omega\cdot m}{3.24\times 10^{-7}\ \ m^2}=4.6\ \ \Omega/m"


(b) Current in wire "I=\dfrac{\Delta V}{R}=\dfrac{10V}{4.6\Omega}=2.2 A"



(3.) "R(150\degree C)=2.415 \Omega=R(0\degree C)(1+\alpha(150-0))\ \ \ ......(i)"

"R(20\degree C)=1.642\ \Omega=R(0\degree C)(1+\alpha(20-0))\ \ \ \ .....(ii)"

Substitute first equation in second and solve for "\alpha"

On Solving, we get

"\alpha=3.9\times 10^{-3}\ \degree C^{-1}"

and "R(0\degree C)=1.5236\ \ \Omega"


With Ξ± known, use any of the above equations to get

"R(0\degree C)=1.5236Ω" . Then use "R(0oC)=ρ(0\degree C)L/A"

"1.5236Ω=\dfrac{\rho(0\degree C)\cdot 300m}{0.25\pi(2\times 10^{-3} m)^2}\Rightarrow \rho(0\degree C)=1.596\times 10^{-8}\ \Omega m"


and "\rho(20\degree C)=\rho(0\degree C)(1+\alpha\cdot 20)=1.72\times 10^{-8}\ \Omega m"